Show that z = i is a root of the equation z
4 + z
3 + z − 1 = 0. Find the three
other roots.
1
Expert's answer
2020-05-20T19:40:03-0400

"z^4 + z^3 + z \u2212 1 = 0"

"z(z^2+1)+z^4-1=0"

"z(z^2+1)+(z^2-1)(z^2+1)=0"

"(z^2+1)(z+z^2-1)=0"

"z^2+1=0, z+z^2-1=0"


"z^2+1=0"

"z^2=-1"

"z_1=i" "z_2=-i"


"z+z^2-1=0"

"D=1+4=5"

"z_3=\frac{-1+\sqrt{5}}{2}" "z_4=\frac{-1-\sqrt{5}}{2}"


Answer

The three other roots are "z_2,z_3,z_4" .